on proportions and percents how would you work the question. what is 5/4% of ___ is 6
Answer by stanbon(75887)

what is 5/4% of ___ is 6
------
Equation:
(5/4)%*x = 6
---
(5/400)x = 6
---
(1/80)x = 6
--
x = 480
==================
cheers,
Stan H.